I need some help with this.
Evaluate.
5 SQRT of 2^5

Answer by stanbon(75887)

5 SQRT of 2^5
5sqrt(2^5)= 5sqrt(32)=5sqrt(16*2)=5(4)sqrt2=20sqrt(2)
Cheers,
Stan H.